# Inductive Meta-path Learning for Schema-complex Heterogeneous Information Networks

### v1

@@ 4.1.3 Comparing the Multi-relation Inductive and Multi-relation Transductive Capacity of SchemaWalk
Figure 4 shows the performance drop of SchemaWalk from multi-relation inductive experiments to multi-relation transductive ones. Impressively, the performance degradation is negligible across all metrics, with an average decline rate of merely $3.32\%$. This observation suggests that SchemaWalk can effectively learn a policy to establish high-quality meta-paths by examining meta-path samples (and their scores) generated during training on some relations. With this policy, the model can efficiently predict meta-paths for test relations by referring to the trained schema-level representations.
@@ 4.2 Per-relation Experiments
Baselines. We compare SchemaWalk with eight existing relation reasoning methods: MPDRL [37], PCRW [38], Autopath [18], Metapath2Vec [30], HIN2Vec [11], RotatE [29], TransE [27], MINERVA [20]. See Appendix A for their detailed information.
@@
Training Settings for SchemaWalk. On the schema graph, a relation could connect multiple entity type pairs $(t_{src},t_{tgt})$. On each attempt, the agent chooses one such pair to form a query, starts at $t_{src}$ and finds a rewarding path to $t_{tgt}$. For YAGO26K-906 and Chem2Bio2RDF, the query set for training is all possible entity type pairs. However, in NELL, some relations could connect an enormous number of entity type pairs (e.g., 1488 for CompetesWith) and many of these type pairs map to very few entity pairs. Therefore, for SchemaWalk, we narrow the query set on NELL by the minimum type pairs that cover a threshold percent (set as $80\%$) of the $r_{q}$-related entity pairs. See Appendix C for further training settings.
@@
SchemaWalk
MPDRL
PCRW
Autopath
Metapath2Vec
HIN2Vec
RotatE
TransE
MINERVA
YAGO26K-906
isCitizenOf ROC-AUC
$\boldsymbol{0.865}$
$0.778$
$0.584$
$0.755$
$0.652$
$0.800$
$0.778$
$0.590$
$0.828$
AP
$\boldsymbol{0.904}$
$0.793$
$0.706$
$0.723$
$0.781$
$0.837$
$0.830$
$0.810$
$0.840$
DiedIn ROC-AUC
$\boldsymbol{0.947}$
$0.710$
$0.645$
$0.723$
$0.661$
$0.785$
$0.860$
$0.622$
$0.632$
AP
$\boldsymbol{0.971}$
$0.679$
$0.686$
$0.787$
$0.830$
$0.877$
$0.907$
$0.745$
$0.786$
GraduatedFrom ROC-AUC
$\boldsymbol{0.896}$
$0.664$
$0.586$
$0.724$
$0.661$
$0.803$
$0.830$
$0.662$
$0.609$
AP
$\boldsymbol{0.912}$
$0.743$
$0.664$
$0.718$
$0.783$
$0.842$
$0.855$
$0.750$
$0.718$
NELL
WorksFor ROC-AUC
$\boldsymbol{0.913}$
$0.759$
$0.646$
$0.703$
$0.613$
$0.790$
$0.868$
$0.637$
$0.767$
AP
$\boldsymbol{0.945}$
$0.871$
$0.735$
$0.778$
$0.819$
$0.870$
$0.911$
$0.719$
$0.802$
PlaysAgainst ROC-AUC $0.907$
$0.774$
$0.567$
$0.544$
$0.761$
$0.783$
$\boldsymbol{0.966}$
$0.845$
$0.541$
AP $0.938$
$0.823$
$0.745$
$0.691$
$0.904$
$0.917$
$\boldsymbol{0.993}$
$0.928$
$0.683$
CompetesWith ROC-AUC
$0.706$
$0.589$
$0.547$
$0.567$
$0.600$
$0.730$
$\boldsymbol{0.870}$
0.7740.774{0.774}0.774 $0.585$
AP
$0.806$
$0.695$
$0.699$
$0.685$
$0.735$
$0.867$
$\boldsymbol{0.939}$
$0.908$ $0.702$
Chem2Bio2RDF
Bind ROC-AUC
$\boldsymbol{1.000}$
$0.813$
$0.609$
$0.685$
$0.759$
$0.880$
$0.936$
$0.909$
$0.715$
AP
$\boldsymbol{1.000}$
$0.890$
$0.723$
$0.792$
$0.914$
$0.939$
$0.984$
$0.978$
$0.816$
@@ 4.2.1 Results on Schema-complex HINs
Among the meta-path based baselines, HIN2Vec achieves decent prediction accuracy. It generates superior embeddings compared to Metapath2Vec, by utilizing meta-paths for multiple relations. MPDRL produces relatively satisfactory results, thanks to its RL-based strategy to explore instance paths. Autopath exhibits acceptable capability, but its ability is confined by the limited number of discovered meta-paths, particularly for relations that involve an extensive meta-path space. PCRW is the most mediocre model among them because of its randomness-based nature. For the multi-hop-based MINERVA, it is performs well in reasoning isCitizenOf and WorksFor, but performs poorly for other relations. We observe that embedding-based approach for KBs, particularly RotatE, dominates the two relations in NELL. Upon further inspection, we notice that both PlaysAgainst and CompetesWith involve thousands of type pair, suggesting an enormous meta-path space (ten million-scale) to explore.
@@ 4.2.3 Entity-level Inductive Experiment
In the Entity-level inductive experiment, we concentrate on the NELL dataset and compared our method against RotatE. We randomly sample $40\%$ of the positive testing set, remove $0\%$, $20\%$, $50\%$ and $100\%$ of the nodes that appears in the pairs from the instance graph, followed by regular training and link prediction. The ROC-AUC and AP results for the two methods under the four scenarios are shown in Fig. 5.
@@
5 Further Analysis
5.1 Convergence Properties
During the initial training period, SchemaWalk exhibits noticeable exploratory behaviour, thanks to the regularization item. Nevertheless, this behaviour does not adversely affect the learning of the agent, as the confidence and coverage curves converges eventually, even for relations PlaysAgainst and CompetesWith that connect thousands of entity type pairs. The arrival rate curves tend to converge faster compared to the confidence and coverage curves, indicating that the agent needs to know how to navigate to the target type node before learning to generate high quality meta-paths. The majority of the confidence/coverage curves end up converging at different values, reflecting the variations in the confidence/coverage distribution for meta-paths amongst different relations. Lastly, all curves converge before the set number of iterations and further training iterations do not empirically yield better prediction results.
@@ 5.2 Meta-paths Analysis
We examine the generated meta-paths and select those meta-paths with high rewards, which are presented in Table VI. All these meta-paths accurately conveyed the semantic meaning of the given relations. We also observe that our model is capable of using synonymous relations to explain target relations, as demonstrated by utilizing CompetesWith for PlayAgainst and Nationality for StateOfOrigin). In multi-relation inductive settings, SchemaWalk also infer high-quality meta-paths for untrained relations (e.g., StateOfOrigin and MusicalBand in DB [^1] K-174).
Besides, our model can mine meta-paths not only for top-hierarchy types like Person, but also for types from lower hierarchies like CEO and Scientist. Their coverage and confidence are evaluated under the relevant KB. We thus observe that 16.7% of the scientists entailing isCitizenOf would have information of where their alma maters locate and, 54.0% of the target countries indicated by such meta-path are the true answers. We could also observe some interesting facts for politicians, scientists and coaches. For example, we find that 12.5% of the scientists would graduate from the same university where their scientist spouse works.
@@ 6 Conclusion and Future Work
Future work includes at least three aspects. First, we could combine SchemaWalk’s rewards with the evaluation metrics for downstream tasks and obtain more task-specific meta-paths. Besides, advances in graph representation learning and RL may enable us to design better agent and further improve the performance. Lastly, we could build a score estimator to reduce the time to evaluate each meta-path during learning.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (NSFC, 62206303, 62001495) and China Postdoctoral Science Foundation (Grant 48919). We would like to express our gratitude to Ziniu Hu for his valuable advice. We would also like to thank Jie Kang and Yuehang Si for providing us with some baseline results.
@@ Appendix A Baseline Implementations
  • RotatE [29]. RotatE555github.com/DeepGraphLearning/KnowledgeGraphEmbedding learns embeddings to represent entities and relations in KBs. We adopt the reported hyperparameters for Yago and apply the same for other datasets.
  • TransE [27]. TransE builds the embeddings of instance triples by making the sum of the head/relation vector close to the tail vector. We use an implementation by Han et al.666github.com/thunlp/OpenKE.
[…]
  • ComplEx [35]. ComplEx migrates DistMult in a complex space and offers comparable performance777github.com/ttrouill/complex.
  • RNNLogic [36]. RNNLogic is a rule-learning method that consists of a recurrent neural network to encode KB into a low-dimensional representation and a logic reasoning module to learn and infer logic rules from the encoded representation888github.com/DeepGraphLearning/RNNLogic.
  • MINERVA [20]. MINERVA is a neural RL-based multi-hop approach for automated reasoning999github.com/shehzaadzd/MINERVA. For per-relation experiments, similar to their evaluation approach on NELL that uses the predicted scores (logits) for head/tail entity pairs, we calculate the predicted scores for all positive and negative samples and obtain the similarity with a softmax operation over these scores, which produces better result than using the original scores.
@@ Appendix B Dataset Details
  • YAGO26K-906 [39, 40]. Yago is a KB built on extracted facts from Wikipedia and WordNet. The original Yago poses a limitation on the semantic relations among entity types. We adopt the preprocessed core Yago facts with enriched taxonomy, released by Hao et al.101010github.com/JunhengH/joie-kdd19.
[…]
@@
  • Chem2Bio2RDF [42]. Chem2Bio2RDF is a schema-simple HIN linking drug candidates and their biological annotations for drug-target prediction111111chem2bio2rdf.org/.
@@
Base Iterations $I_{base}$ 500 (YAGO26K-906)
Base iterations for training
@@
Rounds $I_{r}$ 5
Training rounds for each relation
Meta-path Length $l$ 5
Maximum length for meta-paths
Hidden Size $d_{h}$ 200 (YAGO26K-906)
Dimension of hidden layers output
@@
Embedding Size $d_{e}$ 64
Dimension of embeddings for concepts and relations representation
Batch Size $K$ 20 /80 (YAGO26K-906)
Batch size to sample concept pairs to form queries (batch size for multi-relation / per-relation if two values are given)
@@
Rollouts $N$ 40
Number of rollouts for each training sample
Test Batch Size 10 (YAGO26K-906)
Batch size during test
@@
Test Rollouts 10 (YAGO26K-906)
Beam Search Width
@@
Learning Rate $\alpha$ 0.0005
Update rate for Adam optimizer
Beta $\beta$ 0.05 (decay rate 0.9)
Coefficient for entropy regularization term
Lambdas $\lambda_{1},\lambda_{2}$ 1, 1
Weight in Eq.4
@@
Appendix C Further Training Settings

### v2

@@ 4.1.3 Comparing the Multi-relation Inductive and Multi-relation Transductive Capacity of SchemaWalk
Fig. 4 shows the performance drop of SchemaWalk from multi-relation inductive experiments to multi-relation transductive ones. Impressively, the performance degradation is negligible across all metrics, with an average decline rate of merely $3.32\%$. This observation suggests that SchemaWalk can effectively learn a policy to establish high-quality meta-paths by examining meta-path samples (and their scores) generated during training on some relations. With this policy, the model can efficiently predict meta-paths for test relations by referring to the trained schema-level representations.
@@ 4.2 Per-relation Experiments
Baselines. We compare SchemaWalk with eight existing relation reasoning methods: MPDRL [44], PCRW [45], Autopath [18], Metapath2Vec [31], HIN2Vec [11], NSHE [46], RotatE [30], TransE [28], MINERVA [20]. See Appendix A for their detailed information.
@@
SchemaWalk MPDRL PCRW Autopath Metapath2Vec HIN2Vec NSHE RotatE TransE MINERVA
YAGO26K-906
isCitizenOf ROC-AUC $\boldsymbol{0.865}$ $0.778$ $0.584$ $0.755$ $0.652$ $0.800$ 0.788 $0.778$ $0.590$ $0.828$
AP $\boldsymbol{0.904}$ $0.793$ $0.706$ $0.723$ $0.781$ $0.837$ 0.803 $0.830$ $0.810$ $0.840$
DiedIn ROC-AUC $\boldsymbol{0.947}$ $0.710$ $0.645$ $0.723$ $0.661$ $0.785$ 0.748 $0.860$ $0.622$ $0.632$
AP $\boldsymbol{0.971}$ $0.679$ $0.686$ $0.787$ $0.830$ $0.877$ 0.778 $0.907$ $0.745$ $0.786$
GraduatedFrom ROC-AUC $\boldsymbol{0.896}$ $0.664$ $0.586$ $0.724$ $0.661$ $0.803$ 0.815 $0.830$ $0.662$ $0.609$
AP $\boldsymbol{0.912}$ $0.743$ $0.664$ $0.718$ $0.783$ $0.842$ 0.839 $0.855$ $0.750$ $0.718$
NELL
WorksFor ROC-AUC $\boldsymbol{0.913}$ $0.759$ $0.646$ $0.703$ $0.613$ $0.790$ 0.774 $0.868$ $0.637$ $0.767$
AP $\boldsymbol{0.945}$ $0.871$ $0.735$ $0.778$ $0.819$ $0.870$ 0.859 $0.911$ $0.719$ $0.802$
PlaysAgainst ROC-AUC $0.907$ $0.774$ $0.567$ $0.544$ $0.761$ $0.783$ 0.814 $\boldsymbol{0.966}$ $0.845$ $0.541$
AP $0.938$ $0.823$ $0.745$ $0.691$ $0.904$ $0.917$ 0.883 $\boldsymbol{0.993}$ $0.928$ $0.683$
CompetesWith ROC-AUC $0.706$ $0.589$ $0.547$ $0.567$ $0.600$ $0.730$ 0.677 $\boldsymbol{0.870}$ 0.7740.774{0.774}0.774 $0.585$
AP $0.806$ $0.695$ $0.699$ $0.685$ $0.735$ $0.867$ 0.784 $\boldsymbol{0.939}$ $0.908$ $0.702$
Chem2Bio2RDF
Bind ROC-AUC $\boldsymbol{1.000}$ $0.813$ $0.609$ $0.685$ $0.759$ $0.880$ 0.901 $0.936$ $0.909$ $0.715$
AP $\boldsymbol{1.000}$ $0.890$ $0.723$ $0.792$ $0.914$ $0.939$ 0.963 $0.984$ $0.978$ $0.816$
@@
Training Settings for SchemaWalk. On the schema graph, a relation could connect multiple entity type pairs $(t_{src},t_{tgt})$. On each attempt, the agent chooses one such pair to form a query, starts at $t_{src}$ and finds a rewarding path to $t_{tgt}$. For YAGO26K-906 and Chem2Bio2RDF, the query set for training is all possible entity type pairs. However, in NELL, some relations could connect an enormous number of entity type pairs (e.g., 1488 for CompetesWith) and many of these type pairs map to very few entity pairs. Therefore, for SchemaWalk, we narrow the query set on NELL by the minimum type pairs that cover a threshold percent (set as $80\%$) of the $r_{q}$-related entity pairs. See Appendix D for further training settings.
@@ 4.2.1 Results on Schema-complex HINs
Among the meta-path based baselines, HIN2Vec achieves decent prediction accuracy. It generates superior embeddings compared to Metapath2Vec, by utilizing meta-paths for multiple relations. The performance of NSHE is also reasonably good, similar to HIN2Vec, due to the high-order structure-derived embeddings it obtains. MPDRL produces relatively satisfactory results, thanks to its RL-based strategy to explore instance paths. Autopath exhibits acceptable capability, but its ability is confined by the limited number of discovered meta-paths, particularly for relations that involve an extensive meta-path space. PCRW is the most mediocre model among them because of its randomness-based nature. For the multi-hop-based MINERVA, it performs well in reasoning isCitizenOf and WorksFor, but performs poorly for other relations. We observe that embedding-based approach for KBs, particularly RotatE, dominates the two relations in NELL. Upon further inspection, we notice that both PlaysAgainst and CompetesWith involve thousands of type pair, suggesting an enormous meta-path space (ten million-scale) to explore.
@@ 4.2.3 Entity-level Inductive Experiment
In the Entity-level inductive experiment, we concentrate on the NELL dataset and compared our method against RotatE. We randomly sample $40\%$ of the positive testing set, remove $0\%$, $20\%$, $50\%$ and $100\%$ of the nodes that appear in the pairs from the instance graph, followed by regular training and link prediction. The ROC-AUC and AP results for the two methods under the four scenarios are shown in Fig. 5.
@@
5 Further Analysis
5.1 Convergence Properties
During the initial training period, SchemaWalk exhibits noticeable exploratory behaviour, thanks to the regularization item. Nevertheless, this behaviour does not adversely affect the learning of the agent, as the confidence and coverage curves converge eventually, even for relations PlaysAgainst and CompetesWith that connect thousands of entity type pairs. The arrival rate curves tend to converge faster compared to the confidence and coverage curves, indicating that the agent needs to know how to navigate to the target type node before learning to generate high quality meta-paths. The majority of the confidence/coverage curves end up converging at different values, reflecting the variations in the confidence/coverage distribution for meta-paths amongst different relations. Lastly, all curves converge before the set number of iterations and further training iterations do not empirically yield better prediction results.
With aforementioned notations, the overall complexity for the training process is bounded by $\mathcal{O}(KN(l-1+(|V|\Delta(\mathcal{G}_{I})^{2})^{l-2})$ per epoch (See Appendix E.1 for detailed analysis). Even when training on the most complex schema graph, that of NELL, our method demonstrates remarkable efficiency. The observed wall times for our model to complete a single epoch for the relations PlaysAgainst, CompetesWith, and PlaysFor are 1.44s, 1.98s, and 5.01s respectively. Consequently, it took only about 868s, 793s, and 1353s respectively to converge to a reasonable point for these relations. This observation affirms that our strategy, which involves searching for the target type node within the schema graph, delivers efficient training.
@@ 5.2 Meta-paths Analysis
We examine the generated meta-paths and select those meta-paths with high rewards, which are presented in Table VI. A careful observation of these meta-paths reveals their significant diversity and contextual relevance. For instance, our model can learn multiple meta-paths to deduce citizenship via birthplace, alma mater, or workplace. Furthermore, it learns meta-paths not just for top-hierarchy types like Person, but also for specific types from lower hierarchies like CEO and Scientist. In terms of semantic correctness, each of these meta-paths precisely encapsulates and expresses the semantic meaning of the corresponding relations. An example includes our model learning the meta-path that infers an individual’s alma mater by taking into account the employment location of their university instructor. We also observe that our model is capable of using synonymous relations to explain target relations, as demonstrated by utilizing CompetesWith for PlayAgainst and Nationality for StateOfOrigin). In multi-relation inductive settings, SchemaWalk also infer high-quality meta-paths for untrained relations (e.g., StateOfOrigin and MusicalBand in DB [^1] K-174).
The coverage and confidence of above meta-paths are evaluated under the relevant KB. We thus observe that 16.7% of the scientists entailing isCitizenOf would have information of where their alma maters locate and, 54.0% of the target countries indicated by such meta-path are the true answers. We could also observe some interesting facts for politicians, scientists and coaches. For example, we find that 12.5% of the scientists would graduate from the same university where their scientist spouse works.
5.3 Parameter Analysis on $\lambda_{1}$ and $\lambda_{2}$
Figure 7: The left and right plots respectively depict the results for transductive KG reasoning on DB [^1] K-174 (measured by MRR) and per-relation link prediction on YAGO26K-906 with the relation isCitizenOf (expressed in AUC) with varying values for $(\lambda_{1},\lambda_{2})$.
We scrutinize the impact of parameters $(\lambda_{1},\lambda_{2})$, which signify respective weights for coverage and confidence as expressed in Equation (4), on the performance of our method. We selected values for $\lambda_{1}$ from the set $\{1,5,10,20\}$ and for $\lambda_{2}$ from $\{1,2,5,10\}$. The scale for these parameters is not identical due to a typically smaller value of coverage in relation to confidence. With these selected parameters, we conducted a series of experiments for transductive KG reasoning on DB [^1] K-174 and per-relation link prediction on the YAGO26K-906 (on relation isCitizenOf). Our empirical observations indicate that the optimal values for $(\lambda_{1},\lambda_{2})$ for these distinct tasks are $(10,1)$ and $(5,1)$, respectively. Moreover, it is noteworthy that performance deteriorates by 12% and 7% respectively when $\lambda_{1}$ is excessively large. This downturn in performance is possibly due to the arrival item, devised to facilitate training, being swallowed up in the reward signal when in these extreme cases. However, interestingly, even when faced with overly large values for $(\lambda_{1},\lambda_{2})$, the degradation in performance was not significant. This, in turn, underscores that our approach maintains a smooth performance trajectory irrespective of the values of $(\lambda_{1},\lambda_{2})$, denoting a low sensitivity to these parameters.
5.4 Inference Time
We pose that our model offers superior efficiency during inference since it mainly search for the target entity type within its corresponding local neighborhood indicated by the schema graph. This sharply contrasts with the markedly lower efficiency of other meta-path discovery methods, which necessitate exhaustive exploration of the instance graph. The primary computational expense during inference for SchemaWalk rests in the computation of probabilities for all potential outgoing edges along the defined meta-path. Therefore, the inference time of SchemaWalk is fundamentally contingent on the degree distribution of the graph. Assuming the schema graph $\mathcal{G}_{s}$ conforms to a power law degree distribution, the test-time complexity per epoch, based on the paper’s notations, is given as $\mathcal{O}(B^{\prime}(l-1)\frac{\eta}{\eta-1})$, represents the power law’s coefficient. Please refer to Appendix E.2 for an in-depth analysis.
@@ 6 Conclusion and Future Work
Future work includes at least four aspects. First, we could combine SchemaWalk’s rewards with the evaluation metrics for downstream tasks and obtain more task-specific meta-paths. Besides, advances in graph representation learning and RL may enable us to design better agent and further improve the performance. Furthermore, we could build a score estimator to reduce the time to evaluate each meta-path during learning. Lastly, we could integrate temporal logic into our reasoning method, thereby making it applicable to temporal HIN.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 62206303 and Grant 62001495 and in part by China Postdoctoral Science Foundation under Grant 48919 and Science and Technology Innovation Program of Hunan Province (Grant Number: 2023RC3009). We would like to express our gratitude to Ziniu Hu for his valuable advice. We would also like to thank Jie Kang and Yuehang Si for providing us with some baseline results.
@@ Appendix A Baseline Implementations
  • NSHE [46]. NSHE is a network schema preserving HIN embedding method that retains high-order structure.555github.com/Andy-Border/NSHE
  • RotatE [30]. RotatE666github.com/DeepGraphLearning/KnowledgeGraphEmbedding learns embeddings to represent entities and relations in KBs. We adopt the reported hyperparameters for Yago and apply the same for other datasets.
  • TransE [28]. TransE builds the embeddings of instance triples by making the sum of the head/relation vector close to the tail vector. We use an implementation by Han et al.777github.com/thunlp/OpenKE.
  • DistMult [40]. DistMult associates related entities using Hadamard product of embeddings5.
  • ComplEx [41]. ComplEx migrates DistMult in a complex space and offers comparable performance888github.com/ttrouill/complex.
  • RNNLogic [42]. RNNLogic is a rule-learning method that consists of a recurrent neural network to encode KB into a low-dimensional representation and a logic reasoning module to learn and infer logic rules from the encoded representation999github.com/DeepGraphLearning/RNNLogic.
  • MINERVA [20]. MINERVA is a neural RL-based multi-hop approach for automated reasoning101010github.com/shehzaadzd/MINERVA. For per-relation experiments, similar to their evaluation approach on NELL that uses the predicted scores (logits) for head/tail entity pairs, we calculate the predicted scores for all positive and negative samples and obtain the similarity with a softmax operation over these scores, which produces better result than using the original scores.
  • MLN4KB [43]. MLN4KB is a method based on Markov logic network that can leverage the sparsity of knowledge bases111111github.com/baidu-research/MLN4KB.
@@ Appendix B Dataset Details
  • YAGO26K-906 [47, 48]. Yago is a KB built on extracted facts from Wikipedia and WordNet. The original Yago poses a limitation on the semantic relations among entity types. We adopt the preprocessed core Yago facts with enriched taxonomy, released by Hao et al.121212github.com/JunhengH/joie-kdd19.
  • DB [^1] K-174 [1]. The original Dbpedia is a large KB extracted from Wikipedia involving considerable specific domains and general knowledge. Similar with YAGO26K-906, we use the taxonomy-enriched version10.
@@
  • Chem2Bio2RDF [50]. Chem2Bio2RDF is a schema-simple HIN linking drug candidates and their biological annotations for drug-target prediction131313chem2bio2rdf.org/.
@@
Base Iterations $I_{base}$ 500 (YAGO26K-906) Base iterations for training
@@
Rounds $I_{r}$ 5 Training rounds for each relation
Meta-path Length $l$ 5 Maximum length for meta-paths
Hidden Size $d_{h}$ 200 (YAGO26K-906) Dimension of hidden layers output
@@
Embedding Size $d_{e}$ 64 Dimension of embeddings for concepts and relations representation
Batch Size $K$ 20 /80 (YAGO26K-906) Batch size to sample concept pairs to form queries (batch size for multi-relation / per-relation if two values are given)
@@
Rollouts $N$ 40 Number of rollouts for each training sample
Test Batch Size 10 (YAGO26K-906) Batch size during test
@@
Test Rollouts 10 (YAGO26K-906) Beam Search Width
@@
Learning Rate $\alpha$ 0.0005 Update rate for Adam optimizer
Beta $\beta$ 0.05 (decay rate 0.9) Coefficient for entropy regularization term
Lambdas $\lambda_{1},\lambda_{2}$ 10, 1 (Multi-Relation) Weight in Eq.4
5, 1 (Per-Relation)
@@
Appendix C Supporting Node Classification
TABLE IX: Node classification task on ACM and DBLP.
ACM DBLP
Marco-F1 Mirco-F1 Marco-F1 Mirco-F1
DeepWalk [51] 84.17 83.99 84.73 85.22
Metapath2Vec [31] 73.83 73.67 91.92 92.73
RGCN [52] 91.56 91.37 91.58 92.03
GAT [53] 87.57 87.38 90.79 91.85
HAN [54] 90.93 90.78 91.56 91.96
MAGNN [55] 90.83 90.75 92.93 93.51
HGT [56] 91.18 90.98 92.89 93.46
SchemaWalk-HAN 92.48 92.34 93.26 93.68
SchemaWalk-MAGNN 92.96 92.82 93.37 93.83
While our primary focus is on relation prediction, our method could be combined with existing HIN embedding methods for performing node classification. Specifically, we employ SchemaWalk to identify long meta-path, which are subsequently used to embed HIN with HAN [54] and MAGNN [55]. We follow the basic node classification task as specified in [54] and we use 30% of the target type nodes for training and 70% for testing. As evidenced in Table IX, our method effectively identifies lengthy meta-paths, which when used in conjunction with HAN and MAGNN, yield favourable performance for node classification.
Appendix D Further Training Settings